# Jon Danielsson London School of Economics Andreas Uthemann

### v2

@@ 3.2 Wrong-Way Risk
Wrong-way risk refers to situations where exposure to risk increases when that risk is most severe, and AI introduces particularly dangerous forms of it, stemming from how AI gains trust. AI builds trust incrementally by excelling at simple, repetitive tasks leveraging its information processing advantages. As AI systems demonstrate competence in increasingly complex tasks, financial institutions progressively entrust them with more consequential decisions. This may culminate in an AI version of the Peter principle, where AI systems are promoted to increasingly critical roles until their capabilities no longer match the job requirements.
@@
The most serious wrong-way risk arises from a fundamental mismatch between AI’s information processing strengths and the nature of financial crises. AI excels at pattern recognition based on historical data, but struggles with unprecedented scenarios — the unknown unknowns that characterise financial crises. As Danielsson (2022) argues, every financial crisis is unique in its specifics. While the same three fundamental vulnerabilities — excessive leverage, liquidity preference, and system opacity — cause all crises, the precise mechanisms differ significantly from one crisis to the next. Consequently, while AI can be a considerable aid in identifying stress and analysing crises ex-post (Hellwig, 2021; Fouliard et al., 2021; Batsuuri et al., 2024), it will not be very helpful in predicting and preventing crises.
The problem is exacerbated by the fact that we cannot predict how future financial institutions and regulators will react to stress. Their reaction functions remain unknown because policy makers like to keep strategic ambiguity, rotation of senior decision makers, the legal and political environment will be different, and they do not even know how they will react without knowing the concrete circumstances. Consequently, these reaction functions cannot be captured in historical datasets. Something that is not in the dataset cannot be learned by neural networks, no matter how sophisticated.
@@ 3.3 Procyclicality and Market Structure
AI deployment in financial markets creates a powerful drive to more synchronised behaviour through two related channels: homogenisation of risk assessment methodologies and consolidation of market structure. Both channels reduce diversity in financial decision-making — the first by promoting similar analytical approaches across institutions regardless of size, and the second by concentrating market power among fewer, larger players. This twin reduction in diversity significantly alters how markets absorb and respond to shocks, inducing procyclicality, with important implications for financial stability. Technically, this can happen when the correlation of institutions’ signals rises, so the stabilising heterogeneity in equilibrium behaviour is reduced, increasing the risk of procyclical herd behaviour.
@@ 3.3.2 Market Concentration
The cost structure of AI development further concentrates decision-making. As Gambacorta and Shreeti (2025) observe, developing state-of-the-art AI systems requires significant investment in computing infrastructure, specialised talent, and proprietary data — resources available primarily to the largest financial institutions. This creates a tiered market structure where Globally Systemically Important Banks (GSIBs) can afford to develop proprietary AI engines, giving them significant competitive advantages in risk management and trading, while mid-tier institutions with legacy systems and traditional staff struggle to compete and must rely on off-the-shelf AI solutions from a limited number of vendors. Meanwhile, nimble fintech startups with modern technology stacks can leverage commodity AI tools creatively but lack the resources for fully customised solutions.
@@ 3.4 Speed
The most immediate damage from financial crises stems from financial institutions’ instinct to protect themselves during stress periods. When shocks occur, institutions must rapidly decide whether to run or to stay. This decision depends not only on fundamentals but critically on others’ actions, with severe consequences for wrong decisions.
@@ 4 Optimal Policy Response
AI poses both opportunities and challenges for financial stability, and the authorities must respond to AI. If they do so effectively, they can leverage it to perform their mandate better. When they do not, they risk destabilising the financial system. We outline three critical areas, arising from the model in Section 2 and the vulnerability channels in the previous Section, where financial authorities will have to respond to AI.
@@ 4.1 Authority AI Systems
To begin with, the authorities can benefit from developing their own AI capabilities to design regulations and evaluate intervention effectiveness. When AI first appeared on central banks’ radar, many centred their response around divisions dealing with data, IT, or innovation. However, given how AI can act as a powerful agent for systemic risk, financial stability should play a central role in AI policy. Model link: building in-house authority AI raises the effective cost of a hostile signal shift, reducing the probability that malicious information pushes beliefs past the run threshold.
For authority AI systems to be effective, the data they train on is particularly important. Current approaches to regulation rely on private sector firms providing voluminous PDF reports and database dumps, augmented by supervisors talking to private sector staff. Such a setup does not easily allow authorities to gauge how the private sector would react to potential regulations and interventions. AI can significantly enhance this interaction by opening a new dimension in authority-private sector communication, making the supervisory process more robust and efficient. This requires developing a communication framework — AI-to-AI links — that allows an authority’s AI engine to directly communicate with those of other authorities and the private sector.
@@ 4.2 AI Monitoring Frameworks
Traditional oversight mechanisms focus on capital, liquidity, and conduct, but likely miss the unique systemic risks posed by AI. We propose developing comprehensive AI monitoring frameworks that track how AI is used across the financial system and assess potential coordination risks. To create an effective AI monitoring framework, financial authorities can establish a standardised reporting system — an AI registry — where financial institutions must document all AI systems deployed across their critical functions.
@@ 4.3 Crisis Facilities for AI-Driven Crises
Central banks’ standing facilities will likely become much more important as AI use proliferates. Central banks traditionally prefer discretionary facilities. However, such a deliberative approach may prove too slow in an AI-accelerated crisis. To address this temporal mismatch, automatic, pre-committed liquidity facilities that activate without human intervention would be of benefit. This approach has the additional benefit of potentially ruling out bad equilibria in our game-theoretic framework with public signals. If AI knows that central banks will intervene if prices drop by a certain amount, they will not coordinate on strategies that are only profitable if prices drop more. Automatic standing facilities cap the price drop and raise the cut-off, discouraging the all-exit equilibrium that emerges when signals are highly correlated.
@@ 4.4 Public-Private Partnerships
Another promising avenue is federated learning (McMahan et al., 2017), a technique that enables institutions to train machine learning models locally on sensitive or proprietary datasets, without exposing the underlying data. Instead of sharing raw data, only the trained model parameters or updates are transmitted to a central authority or aggregator. This approach enhances privacy and security while still allowing regulators to benefit from insights across multiple institutions. Sharing model architectures and network parameters a limits the private speed advantage and dampens the procyclical rise in price-impact that accelerates crises.
@@ 5 Conclusion
This paper examines the implications of artificial intelligence for financial stability through a game-theoretic framework. We identify three key mechanisms through which AI affects financial markets: superior information processing, common information sources, and speed advantages. These mechanisms interact with endogenous system responses and strategic complementarities to create both stabilising and destabilising effects.
@@
The financial authorities face significant challenges in adapting to this new landscape. Traditional regulatory frameworks designed for human decision-making timeframes may prove inadequate in markets where critical decisions unfold at machine speed. We propose specific policy responses including authority AI systems, AI-to-AI communication frameworks, triggered facilities, public-private partnerships, and enhanced monitoring systems.
Ultimately, whether AI increases or decreases financial stability depends on how effectively both private market participants and public authorities harness its capabilities. If financial authorities engage proactively with AI technology — developing their own capabilities, establishing direct communication channels with private sector AI, and designing appropriate automated response mechanisms — they can leverage AI to enhance financial stability. If they do not, the likelihood of AI-amplified financial crises will increase. The key to financial stability in the age of AI lies not in resisting technological change, but in ensuring that our regulatory frameworks evolve alongside it.
@@ A.2 Deriving the Threshold Equilibrium
    In the special case where humans succeed with probability $p<1$ when choosing to ”run” while AI agents succeed with probability 1, the analysis from Sakovics and Steiner (2012) implies:
@@ A.3 Speed Advantage and Strategic Uncertainty
The speed advantage of AI is modeled following the approach of Sakovics and Steiner (2012) for heterogeneous withdrawal capabilities. In this setup, the critical threshold $\theta^{*}$ increases with the measure $\mu$ of AI agents, making coordination on the “run” equilibrium more likely.

### v3

@@ 3.2 Wrong-Way Risk
Wrong-way risk refers to situations where exposure to risk increases when that risk is most severe. AI introduces particularly dangerous forms of wrong-way risk, stemming from how it gains trust. AI builds trust incrementally by excelling at simple, repetitive tasks, leveraging its information processing advantages. As AI systems demonstrate competence in increasingly complex tasks, financial institutions progressively entrust them with more consequential decisions. This may culminate in an AI version of the Peter principle, where AI systems are promoted to increasingly critical roles until their capabilities no longer match the job requirements.
@@
The most serious wrong-way risk arises from a fundamental mismatch between AI’s information processing strengths and the nature of financial crises. AI excels at pattern recognition based on historical data but struggles with unprecedented scenarios — the unknown unknowns that characterise financial crises. As Danielsson (2022) argues, every financial crisis is unique in its specifics. While the same three fundamental vulnerabilities — excessive leverage, liquidity preference, and system opacity — cause all crises, the precise mechanisms differ significantly from one crisis to the next. Consequently, while AI can be a considerable aid in identifying stress and analysing crises ex-post (Hellwig, 2021; Fouliard et al., 2021; Batsuuri et al., 2024), it will not be very helpful in predicting and preventing crises.
The problem is exacerbated by the fact that we cannot predict how future financial institutions, political leaders and regulators will react to stress. Their reaction functions remain unknown because policymakers like to keep strategic ambiguity, while rotating senior decision-makers. The legal and political environment will be different, and they do not even know how they will react without knowing the concrete circumstances. Consequently, these reaction functions cannot be captured in historical datasets. Something that is not in a dataset cannot be learned by neural networks, no matter how sophisticated.
@@ 3.3 Procyclicality and Market Structure
AI deployment in financial markets creates a powerful drive to more synchronised behaviour through two related channels: homogenisation of risk assessment methodologies and consolidation of market structure. Both channels reduce diversity in financial decision-making — the first by promoting similar analytical approaches across institutions regardless of size and the second by concentrating market power among fewer, larger players. This twin reduction in diversity significantly alters how markets absorb and respond to shocks, inducing procyclicality with important implications for financial stability. Technically, this can happen when the correlation of institutions’ signals rises, so the stabilising heterogeneity in equilibrium behaviour is reduced, increasing the risk of procyclical herd behaviour.
@@ 3.3.2 Market Concentration
The cost structure of AI development further concentrates decision-making. As Gambacorta and Shreeti (2025) observe, developing state-of-the-art AI systems requires significant investment in computing infrastructure, specialised talent and proprietary data — resources available primarily to the largest financial institutions. This creates a tiered market structure where globally systemically important banks (GSIBs) can afford to develop proprietary AI engines, giving them significant competitive advantages in risk management and trading, while mid-tier institutions with legacy systems and traditional staff struggle to compete and must rely on off-the-shelf AI solutions from a limited number of vendors. Meanwhile, nimble fintech startups with modern technology stacks can leverage commodity AI tools creatively but lack the resources for fully customised solutions.
@@ 3.4 Speed
The most immediate damage from financial crises stems from financial institutions’ instinct to protect themselves during stress periods. When shocks occur, institutions must rapidly decide whether to run or to stay. This decision depends not only on fundamentals but critically on the action of others, with severe consequences for wrong decisions.
@@ 4 Optimal Policy Response
AI poses both opportunities and challenges for financial stability, and the authorities must respond to it. If they do so effectively, they can leverage AI to perform their mandate better. When they do not, they risk destabilising the financial system. We outline four critical areas arising from the model in Section 2 and the vulnerability channels in the previous Section, where financial authorities will have to respond to AI.
@@ 4.1 Authority AI Systems
To begin with, the authorities can benefit from developing their own AI capabilities to design regulations and evaluate intervention effectiveness. When AI first appeared on central banks’ radar, many centred their response around divisions dealing with data, IT or innovation. However, given how AI can act as a powerful agent for systemic risk, financial stability should play a central role in AI policy. Model link: building in-house authority AI raises the effective cost of a hostile signal shift, reducing the probability that malicious information pushes beliefs past the run threshold.
For authority AI systems to be effective, the data they train on is particularly important. Current approaches to regulation rely on private sector firms providing voluminous PDF reports and database dumps, augmented by supervisors talking to private sector staff. Such a setup does not easily allow the authorities to gauge how the private sector would react to potential regulations and interventions. AI can significantly enhance this interaction by opening a new dimension in authority-private sector communication, making the supervisory process more robust and efficient. This requires developing a communication framework — AI-to-AI links — that allows an authority’s AI engine to directly communicate with those of other authorities and the private sector.
@@ 4.2 AI Monitoring Frameworks
Traditional oversight mechanisms focus on capital, liquidity, and conduct but likely miss the unique systemic risks posed by AI. We propose developing comprehensive AI monitoring frameworks that track how AI is used across the financial system and assess potential coordination risks. To create an effective AI monitoring framework, financial authorities can establish a standardised reporting system — an AI registry — where financial institutions must document all AI systems deployed across their critical functions.
@@ 4.3 Crisis Facilities for AI-Driven Crises
Central banks’ standing facilities will likely become much more important as AI use proliferates. Central banks traditionally prefer discretionary facilities. However, such a deliberative approach may prove too slow in an AI-accelerated crisis. To address this temporal mismatch, automatic, pre-committed liquidity facilities that activate without human intervention would be of benefit. This approach has the additional benefit of potentially ruling out bad equilibria in our game-theoretic framework with public signals. If AI knows central banks will intervene if prices drop by a certain amount, they will not coordinate on strategies that are only profitable if prices drop more. Automatic standing facilities cap the price drop and raise the cut-off, discouraging the all-exit equilibrium that emerges when public signals are sufficiently precise.
@@ 4.4 Public-Private Partnerships
Another promising avenue is federated learning (McMahan et al., 2017), a technique that enables institutions to train machine learning models locally on sensitive or proprietary datasets without exposing the underlying data. Instead of sharing raw data, only the trained model parameters or updates are transmitted to a central authority or aggregator. This approach enhances privacy and security while still allowing regulators to benefit from insights across multiple institutions. Sharing model architectures and network parameters limits the private speed advantage and dampens the procyclical rise in price impact that accelerates crises.
@@ 5 Conclusion
This paper examines the implications of artificial intelligence for financial stability through a game-theoretic framework. We identify three key mechanisms through which AI affects financial markets: superior information processing, common information sources and speed advantages. These mechanisms interact with endogenous system responses and strategic complementarities to create both stabilising and destabilising effects.
@@
The financial authorities face significant challenges in adapting to this new landscape. Traditional regulatory frameworks designed for human decision-making timeframes may prove inadequate in markets where critical decisions unfold at machine speed. We propose specific policy responses, including authority AI systems, AI-to-AI communication frameworks, triggered facilities, public-private partnerships, and enhanced monitoring systems.
Ultimately, whether AI increases or decreases financial stability depends on how effectively both private market participants and public authorities harness its capabilities. If the financial authorities engage proactively with AI technology — developing their capabilities, establishing direct communication channels with private sector AI, and designing appropriate automated response mechanisms — they can leverage AI to enhance financial stability. If they do not, the likelihood of AI-amplified financial crises will increase. The key to financial stability in the age of AI lies not in resisting technological change but in ensuring that our regulatory frameworks evolve alongside it.
@@ A.2 Deriving the Threshold Equilibrium
    In the special case where humans succeed with probability $p<1$ when choosing to run while AI agents succeed with probability 1, the analysis from Sakovics and Steiner (2012) implies:
@@ A.3 Speed Advantage and Strategic Uncertainty
The speed advantage of AI is modelled following the approach of Sakovics and Steiner (2012) for heterogeneous withdrawal capabilities. In this setup, the critical threshold $\theta^{*}$ increases with the measure $\mu$ of AI agents, making coordination on the run equilibrium more likely.